a supermarket has been selling discounted apples in bundles of five at their counters. a random sample of 49 bundles weighs 970 grams on average, with a standard deviation of 70 grams. test the hypothesis that =1000 grams against the alternative hypothesis of < 1000 at 0.06 level of significance.


1
Expert's answer
2021-01-27T15:14:45-0500

Hypothesis testing for a mean (σ is unknown, and the variable is normally distributed in the population or n > 30 )

The provided sample mean is "\bar{x}=970" and the sample standard deviation is "s=70," and the size of the sample is "n=49."

The following null and alternative hypotheses need to be tested:

"H_0:\mu=1000"

"H_1:\mu<1000"

This corresponds to a left-tailed test, for which a t-test for one mean, with unknown population standard deviation will be used.

The number of degrees of freedom are "df=n-1=49-1=48," and the significance level is "\alpha=0.06." Based on the provided information, the critical t-value for "\alpha=0.06" and "df=49" degrees of freedom is "t_c=-1.582951."

The rejection region for this left-tailed test is "R=\{t:t<-1.582951\}"

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{970-1000}{70/\sqrt{49}}=-3"

Since it is observed that "t=-3<-1.582951=t_c," it is then concluded that the null hypothesis is rejected. Therefore, there is enough evidence to claim that the population mean "\mu" is less than 1000, at the 0.06 significance level.

Use the P-value approach.

The P-value "t=-3, df=49 \alpha=0.06," left-tailed, is "p=0.002136." Since "p=0.002136<0.06=\alpha," it is then concluded that the null hypothesis is rejected. Therefore, there is enough evidence to claim that the population mean "\mu" is less than 1000, at the 0.06 significance level.